LOFAR Status Meetings

LOFAR Status Meetings (LSM), presenting updates of the status of the telescope and discussing recent developments, are held monthly at ASTRON. The meetings are also broadcast via Vidyo to everyone interested.

Please use a headset and please mute the microphone in the Evo session as long as you do not want to raise a question or say something.

Programme and Presentations

LSM Archive2016201520142013201220112010200920082007


12 December 2012
R. Pizzo Observatory update
G. Heald MSSS update
J. Hessels Pulsar Busy Week update
R. Fallows Recent IPS results
A. Deller Correcting international stations dispersive delays with TEC models
R. Oonk Radio Spectroscopy with LOFAR
28 November 2012
E. Orru' Observatory update
H. Munk Software status update
G. Heald MSSS update
S. Yatawatta Recent results on NCP imaging
V. Jelic Recent polarimetric results for the Elais field
14 November 2012
R. Fallows Observatory update
E. Orru' Status of MSSS "automatic" processing
E. Orru' Update from Imaging Busy Week 15
J. Hessels Follow up on full core tied-array observations using the single clock extension
C. Ferrari LOFAR Source Finders Working Group
31 October 2012
M. Brentjens Observatory Update
G. Heald MSSS update
M. Brentjens Update on the deployment of new syncOptic boards at the CS
J. Hessels (Full core) single clock tests
S. Yatawatta Recent NCP EoR observations
17 October 2012
C. Toribio Observatory Update
D. Mulcahy Investigation on Peeling with Extended Emission and HBA Filter sensitivity
S. Thoudam Octocopter measurements
L. Birzan LBA Observations of Hercules A: preliminary results
F. Breitling Update on the Solar Imaging Pipeline & Solar Calibrator Survey
3 October 2012
W. Frieswijk Radio Observatory update
H. Munk Software status update
G. Heald MSSS update
V. Jelic Commissioning results for the Elais field
J. McKean Commissioning results for Cyg A - to be uploaded
12 September 2012
R. Pizzo Observatory update
G. Heald MSSS update
M. Brentjens Update on the installation of the new synoptic boards
O. Wucknitz Effelsberg sensitivity problem
E. Orru' Commissioning results for DDRG 1835+62
A. Shulevsky Calibration approach for LBA data
22 August 2012
R. Pizzo Observatory update
R. Pizzo MSSS update
H. Munk Software update
O. Wucknitz Phasing up the Superterp – notes on calibration and S/N improvement
R. Fallows Musings on Beam Profiles and Scintillation
11 July 2012
F. Breitling Status of the Solar Imaging Pipeline & Solar Data Center
27 June 2012
A. Polatidis Observatory update
G. Heald MSSS update
J. Hessels Update on single clock on the core stations
S. Buitink LBA bandpass from Cosmic Ray observations
S. Wijnholds EM-modeling of LBA response
13 June 2012
A. Polatidis Observatory update
D. Mulcahy Calibration transfer and interpolation n the M51/3C295 observations
N. Ebbendorf Update on Station Rollout
30 May 2012
A. Polatidis Observatory update
J. Hessels (Full Core) Single Clock tests
A. Deller Update on Long Baseline Working Group busy week
R. Fallows IPS Observations: First comparisons
16 May 2012
A. Polatidis Observatory update
H. Munk Software Status Update
J. Hessels Update on Tied-Array Beams
M. Brentjens Update on PPS synchronization on CS401
H. Holties Long Term Archive Update
18 Apr 2012
R. F. Pizzo Observatory update
M. Wise Commissioning proposals
J. Broderick Update on NCP imaging
A. Gunst Update on PPS synchronization
G. Heald Status beam tests
R. Fallows Solar wind observations with LOFAR: recent results
J. Hessels Pulsar Busy Week update
C. Sobey Measuring ionospheric variations using Pulsar polarimetry with LOFAR
G. Heald MSSS update
04 Apr 2012
A. Polatidis Observatory update
M. Brentjens Update on RSP board synchronization
G. Heald MSSS update
21 Mar 2012
R. Pizzo Observatory update
H. Munk Software development update
A. Corstanje TBB stations calibration and system monitoring
M. Brentjens Update on system sensitivity
J. Hessels Update from Pulsar Busy Week
G. Heald MSSS update
07 Mar 2012
A. Polatidis Observatory update
G. de Bruyn Activities of the Ionospheric Group
A. Bonafede A-team subtraction using SAGECal
G. Heald MSSS update
22 Feb 2012
H. Munk Update on software issues
16 Feb 2012
G. de Bruyn Lofar Sensitivity
J. Hessels Sensitivity Considerations for Beam Formed observations
R. Pizzo Practical Processing Times and Achievable Sensitivities
08 Feb 2012
A. Polatidis Observatory Update
H. Munk Update on software issues
G. Heald MSSS update
M. Brentjens HBA Sensitivity Issues
M. Mevius Update on TBB Busy Week
A. Horneffer Update on the Magnetism KSP Busy Week
R. Pizzo Update on Imaging Busy Week
25 Jan 2012
A. Polatidis Observatory Update
R. Vermeulen Initial Call for Proposals
G. deBruyn Results on HBA SEFD's and bandpass (MHz) ripple
A. Horneffer Update on the Magnetism KSP busy Week
S. Yatawatta Preliminary results on station beam fitting
R. Pizzo Update on Imaging Busy Wednesdays
G. Heald MSSS update
11 Jan 2012
A. Polatidis Observatory Update
M. Mevius TBB Busy Week: Status and plans
A. Corstanje Cable Delay Fitting
  • Last modified: 2017-06-07 10:53
  • by Henk Vosmeijer